OpenStudy (anonymous):

what is the vertex of the function y=x^2-4x+3/

OpenStudy (anonymous):

@agentx5 can you help me

OpenStudy (anonymous):

vertex for of this type of equation: \[y=a(x-h)^{2}+k\] where (h,k) is the vertex. So compleat the square to find this form and you got it

OpenStudy (anonymous):

vertex form*

OpenStudy (anonymous):

y=x^2-2.2x+3+1-1 y=x^2-2.2.x+2^2-1 y+1=(x-2)^2 y=(x-2)^2-1 centre is (2,0)

OpenStudy (anonymous):

jasminecrump18 do u understood

OpenStudy (anonymous):

First of all I would take a look at factoring that quadratic polynomial, it turns into: y=(x-3)(x-1) Solve for x when y=0 to find the roots, +3 and +1 in this case:  The standard form is: \(y = ax^2 + bx + c\) \(y = a(x – h)(x – h) + k\) \(y = ax^2 – 2ahx + ah^2 + k\) The vertex form is: \(y = a(x – h)^2 + k\) Where (h,k) is your vertex's coordinates! So for a, b, c... the x-coordinate is \(-\huge \frac{b}{2a}\) To find the y-coordinate you need to plug back in. PS: Grr @myko